### v2

# A Large and Balanced Corpus for Fine-grained Arabic Readability Assessment

Abstract
A Large and Balanced Corpus
for Fine-grained Arabic Readability Assessment
[…]
A Large and Balanced Corpus
for Fine-grained Arabic Readability Assessment
[…]
1 Introduction
Text readability impacts understanding, retention, reading speed, and engagement DuBay (2004). Texts above a student’s readability level can lead to disengagement Klare (1963). Nassiri et al. (2023) highlighted that readability and legibility depend on both external features (e.g., production, fonts) and content. Text leveling in classrooms helps match books to students’ reading levels, promoting independent reading and comprehension Allington et al. (2015). Developing readability models is crucial for improving literacy, language learning, and academic performance. Readability levels have long been a key component of literacy teaching and learning. One of the most widely used systems in English literacy is Fountas and Pinnell Fountas and Pinnell (2006), which employs qualitative measures to classify texts into 27 levels (A to Z+), spanning from kindergarten to adult proficiency. Similarly, Taha-Thomure (2017)’s system for Arabic has 19 levels from Arabic letters \<أ> A to \<ق> Q. These fine-grained levels are designed for pedagogical effectiveness, ensuring young readers experience gradual, measurable progress, particularly in early education (K–6) Barber and Klauda (2020). A key advantage is that they can be easily mapped to coarser levels with fewer categories, which may be more efficient for broader applications in readability research and automated assessments. In this paper we present the Balanced Arabic Readability Evaluation Corpus (Barec) – a large-scale fine-grained readability assessment corpus across a broad space of genres and readability levels. Inspired by the Taha/Arabi21 readability reference Taha-Thomure (2017), which has been instrumental in tagging over 9,000 children’s books, Barec seeks to establish a standardized framework for evaluating sentence-level [^3] We use sentence to refer to any standalone text segment, including phrases and single words (e.g., book titles). Arabic text readability across 19 distinct levels, ranging from kindergarten to postgraduate comprehension. Our contributions are: (a) a large-scale curated corpus with 69K+ sentences (1M+ words) spanning diverse genres; and (b) benchmarking of automatic readability assessment models across multiple granularities, including both fine-grained (19 levels) and collapsed tiered systems (e.g., five-level and three-level scales) to support various research and application needs, aligning with previous Arabic readability frameworks Al Khalil et al. (2018); Al-Khalifa and Al-Ajlan (2010).
2 Related Work
Automatic Readability Assessment
Automatic readability assessment has been widely studied, resulting in numerous datasets and resources Collins-Thompson and Callan (2004); Pitler and Nenkova (2008); Feng et al. (2010); Vajjala and Meurers (2012); Xu et al. (2015); Xia et al. (2016); Nadeem and Ostendorf (2018); Vajjala and Lučić (2018); Deutsch et al. (2020); Lee et al. (2021). Early English datasets were often derived from textbooks, as their graded content naturally aligns with readability assessment Vajjala (2022). However, copyright restrictions and limited digitization have driven researchers to crowdsource readability annotations from online sources Vajjala and Meurers (2012); Vajjala and Lučić (2018) or leverage CEFR-based L2 assessment exams Xia et al. (2016).
Arabic Readability Efforts
Arabic readability research has focused on text leveling and assessment across various frameworks. Taha-Thomure (2017) proposed a 19-level system for children’s books based on qualitative and quantitative criteria. Other efforts applied CEFR leveling to Arabic, including the KELLY project’s frequency-based word lists, manually annotated corpora Habash and Palfreyman (2022); Naous et al. (2024), and vocabulary profiling Soliman and Familiar (2024). El-Haj et al. (2024) introduced DARES, a readability assessment dataset collected from Saudi school materials. The SAMER project Al Khalil et al. (2020) developed a lexicon with a five-level readability scale, leading to the first manually annotated Arabic parallel corpus for text simplification Alhafni et al. (2024). Automated readability assessment has also been explored through rule-based and machine learning approaches. Early models relied on surface-level features like word and sentence length Al-Dawsari (2004); Al-Khalifa and Al-Ajlan (2010), while later work incorporated POS-based and morphological features Forsyth (2014); Saddiki et al. (2018). The OSMAN metric El-Haj and Rayson (2016) leveraged script markers and diacritization, and recent efforts Liberato et al. (2024) achieved strong results using pretrained models on the SAMER corpus. Building on these efforts, we curated the Barec corpus across genres and readability levels, and manually annotated it at the sentence-level based on an adaptation of Taha/Arabi21 guidelines Taha-Thomure (2017), offering finer-grained control and a more objective assessment of textual variation.
Figure 1: The Barec Pyramid illustrates the relationship across Barec levels and linguistic dimensions, three collapsed variants, and education grades.
Table 1: Representative subset of examples of the 19 Barec readability levels, with English translations, and readability level reasoning. Underlining is used to highlight the main keys that determined the level.
3 Barec Corpus Annotation
In this section, we summarize the guidelines and annotation process. For more details, see Habash et al. (2025). In the next section, we discuss corpus selection and statistics.
3.1 Barec Guidelines
We present below a summarized version of the Barec annotation guidelines. A detailed account of the adaptation process from Taha-Thomure (2017)’s guidelines is in Habash et al. (2025).
Readability Levels
The readability level system of Taha-Thomure (2017) uses the Abjad order of Arabic letters for 19 levels: 1-alif, 2-ba, 3-jim, through to 19-qaf. This system emphasizes a finer distinction in the lower levels, where readability is more varied. The Barec pyramid (Figure 1) illustrates the scaffolding of these levels and their mapping to, guidelines components, school grades, and three collapsed versions of level size 7, 5, and 3. All four level types (19-7-5-3) are fully aligned to easy mapping from fine-grained to coarse-grained levels. We present results for these levels in Section 6.
Readability Annotation Principles
The guidelines focus on readability and comprehension, considering the ease of reading and understanding for independent readers. The evaluation does not depend on grammatical analysis or rhetorical depth but rather on understanding basic, literal meanings. Larger texts may contain sentences at different readability levels, but we focus on sentence-level evaluation, ignoring context and author intent.
Textual Features
Levels are assessed in six key dimensions. Each of these specify numerous linguistic phenomena that are needed to qualify for being ranked in a harder level. Annotators assign each sentence a readability level based on its most difficult linguistic phenomenon. The Cheat Sheet used by the annotators in Arabic and its translation in English are included in Appendix A.
  1. 1.
    Spelling: Word length and syllable count affect difficulty.
  2. 2.
    Word Count: The number of unique words determines the highest level for easier levels.
  3. 3.
    Morphology: We distinguish between simple and complex morphological forms including the use of clitics and infrequent inflectional features, such as the dual.
  4. 4.
    Syntax: Specific sentence structure and syntactic relation constructions are identified as pivotal for certain levels.
  5. 5.
    Vocabulary: The complexity of word choices is key, with higher levels introducing more technical and classical literature vocabulary.
  6. 6.
    Content: The required prior knowledge and abstraction levels are considered for higher levels.
The Barec pyramid (Figure 1) illustrates which aspects are used (broadly) for which levels. For example, spelling criteria are only used up to level 7-zay, while syntax is used until level 15-sin, and word count is not used beyond level 11-kaf.
Problems and Difficulties
Annotators are encouraged to report any issues like spelling errors, colloquial language, or problematic topics. Difficulty is noted when annotations cannot be made due to conflicting guidelines. A few representative examples for each level are provided in Table 1. A full set of examples with explanations of leveling choices is in Appendix A.3.
Table 2: Summary statistics of the Barec Corpus.
3.2 Annotation Team and Process
Annotation Team
The Barec annotation team comprised six native Arabic speakers, all of whom are experienced Arabic language educators. Among the team members, one individual (A0) brought prior experience in computational linguistic annotation projects, while the remaining five (A1-5) possessed extensive expertise in readability leveling, gained through their involvement in the Taha/Arabi21 project.
Annotation Process
The annotation process began with A0, who led sentence-level segmentation and initial text flagging and selection. We followed the Arabic sentence segmentation guidelines by Habash et al. (2022a). Subsequently, A1-5 were tasked with assigning readability labels to the individually segmented texts. The annotation was done through a simple Google Sheet interface. A1-5 received folders containing annotation sets, comprising 100 randomly selected sentences each. The average annotation speed was around 2.5 hours per batch (1.5 minutes/sentence). Before starting the annotation, all annotators received rigorous training, including three pilot rounds. These rounds provided opportunities for detailed discussions of the guidelines, helping to identify and address any issues. 19 shared annotation sets (100 sentence each) were included covertly to ensure quality and measure inter-annotator agreement (IAA). Finally, we conducted a thorough second review of the corpus data, resulting in every sentence being checked twice for the first phase (10,658 sentences) before continuing to finish the 69,441 sentences (1M words). In total, the annotators annotated 92.6K sentences, 25% of which is not in the final corpus: 3.3% were deemed problematic (typos and offensive topics); 11.5% were part of the second round of first phase annotation; and 10.3% were part of the IAA efforts, not including their unification. We report on IAA in Section 6.1.
4 Barec Corpus
Corpus Selection
In the process of corpus selection, we aimed to cover a wide educational span as well as different domains and topics. We collected the corpus from 1,922 documents, which we manually categorized into three domains: Arts & Humanities, Social Sciences, and STEM (details in Appendix C.2) and three readership groups: Foundational, Advanced, and Specialized (details in Appendix C.3). Table 2 shows the distribution of the documents, sentences, and words across domains and groups. The distribution across readership levels aligns with the corpus’s educational focus, with a higher-than-usual proportion at foundational levels. Variations across domains reflect differences in the availability of texts and reader interest (more Arts & Humanities, less STEM). The corpus uses documents from 30 different resources. All selected texts are either out of copyright, within the fair-use limit, or obtained in agreement with publishers. The decision of selecting some of these resources is influenced by the fact that other annotations exist for them. Around 25% of all sentences came from completely new sources that were manually typed to make them digitally usable. All details about the resources are available in Appendix C.
Figure 2: The distribution of the readership groups across Barec levels.
Table 3: Barec Corpus splits.
Readability Statistics
Figure 2 shows the distribution of the three readership groups across all readability levels. As expected, foundational texts strictly dominate the lower levels up to 9-ta, then the presence of advanced and specialized texts starts increasing gradually till the highest level. Specialized texts dominate the highest levels, while the middle levels (10-ya to 14-nun) include a mix of the three groups with a slight advantage for advanced texts.
Corpus Splits
We split the corpus into Train ($\simeq$80%), Dev ($\simeq$10%), and Test ($\simeq$10%) at the document level. Sentences from IAA studies are divided between all splits. However, We will release the IAA studies as a special set as they provide multiple references from different annotators for each example.LABEL:barec-site Also, if other annotations exist for a resource (e.g., CamelTB (Habash et al., 2022b) and ReadMe++ (Naous et al., 2024)), we follow the existing splits. Table 3 shows the corpus splits in the level of documents, sentences, and words. More details about the splits across readability levels, domains, and readership groups are available in Appendix B.
5 Experiments
5.1 Metrics
In this paper, we define the task of Readability Assesment as an ordinal classification task. We use the following metrics for evaluation.
Accuracy (Acc19)
The percentage of cases where reference and prediction classes match in the 19-level scheme. We addition consider three variants, Acc7, Acc5, Acc3, that respectively collapse the 19-levels into the 7, 5, and 3-level schemes discussed in Section 3.
Adjacent Accuracy ($\pm$1 Acc19)
Also known as off-by-1 accuracy. It allows some tolerance for predictions that are close to the true labels. It measures the proportion of predictions that are either exactly correct or off by at most one level.
Average Distance (Dist)
Also known as Mean Absolute Error (MAE), it measures the average absolute difference between predicted and true labels.
Quadratic Weighted Kappa (QWK)
An extension of Cohen’s Kappa (Cohen, 1968; Doewes et al., 2023) measuring the agreement between predicted and true labels, but applies a quadratic penalty to larger misclassifications, meaning that predictions farther from the true label are penalized more heavily. We consider Quadratic Weighted Kappa as the primary metric for selecting the best system.
Table 4: Example sentence in different input variants.
5.2 Input Variants
In morphologically rich languages, affixation, compounding, and inflection convey key linguistic information that influences readability. Human annotators consider morphological complexity when assessing readability, but standard tokenization may obscure these cues. Segmenting sentences into morphological units helps preserve structural patterns relevant to readability prediction. We generate four input variants using CamelTools morphological disambiguation to identify top choice analysis in context (Obeid et al., 2020). [^4] CamelTools v1.5.5: Bert-Disambig+calima-msa-s31 db. For the Word variant, we simply tokenize the sentences and remove diacritics and kashida using CAMeL Tools (Obeid et al., 2020). For Lex, we replace each word with its predicted Lemma. For D3Tok, we tokenize the word into its base and clitics form; and for D3Lex, we replace the base form in D3Tok with the lemma. All variants are dediacritized. Table 4 shows an example of a sentence and the corresponding input variants.
5.3 Fine-Tuning
We fine-tuned the top three Arabic BERT-based models according to Inoue et al. (2021) (AraBERTv02 (Antoun et al., 2020), MARBERTv2 (Abdul-Mageed et al., 2021), CamelBERT-msa (Inoue et al., 2021)). We also added AraBERTv2 to our experiments due to the possible matching between its pre-training data (morphologically segmented sentences by Farasa (Darwish and Mubarak, 2016)) and the different input variants.
5.4 Loss Functions
Since readability levels exhibit a natural ordering, we explore loss functions that account for the distance between predicted and true labels (Heilman et al., 2008). In addition to standard cross-entropy loss (CE), we experiment with Ordinal Log Loss (OLL) (Castagnos et al., 2022), Soft Labels Loss (SOFT) (Bertinetto et al., 2020), Earth Mover’s Distance-based loss (EMD) (L. Hou, 2017), and Regression using Mean Squared Error (Reg) as these have been previously used for ordinal classification tasks. OLL, SOFT, and EMD incorporate a distance matrix $D$ into their formulations to penalize predictions proportionally to their distance from the true label. For simplicity, we define the distance between any two adjacent levels as one, setting $D(i,j)=|i-j|$ for labels i and j. For regression, we round the final output to the nearest readability level to ensure predictions align with the 19 levels.
5.5 Hyper-parameters
For all experiments, we use a learning rate of $5\times 10^{-5}$, a batch size of $64$, and train for six epochs on an NVIDIA V100 GPU. After training, we select the best-performing epoch based on evaluation loss. For Ordinal Log Loss (OLL), we experiment with different values of the weighting parameter $\alpha$, choosing from {0.5, 1, 1.5, 2}. Similarly, for Soft Labels Loss (SOFT), we evaluate different values of the smoothing parameter $\beta$, selecting from {2, 3, 4, 5}. The training of the models in this paper took approximately 20 hours.
5.6 Procedure
Our experiments involve three main variables: the pretrained model, the input variant, and the loss function. Our goal is to determine the optimal combination of these three factors. Due to the large number of experiments required, we divide the process into two stages. In Stage 1, we train all combinations of pretrained models and input variants using cross-entropy loss. We then select the best combination based on a majority vote from our primary evaluation metrics (Acc, Acc $\pm$1, Dist, and QWK). In Stage 2, we take the best combination of pretrained model and input variant from the first stage and train models using all the different loss functions.
6 Results
6.1 Inter-Annotator Agreement (IAA)
In this section, we report on 16 IAA studies, excluding the three pilots and first two IAAs, which overlapped with annotator training.
Pairwise Agreement
The average pairwise exact-match over 19 Barec levels between any two annotators is only 61.1%, which reflects the task’s complexity. Allowing a fuzzy match distance of up to one level raises the match to 74.4%. The overall average pairwise level difference is 0.94 levels. The average pairwise Quadratic Weighted Kappa 81.8% (substantial agreement) confirms most disagreements are minor Cohen (1968); Doewes et al. (2023).
Unification Agreement
After each IAA study, the annotators discussed and agreed on a unified readability level for each sentence. On average, the exact match between the annotators and the unified level (Acc19) was 71.7%, reflecting the difficulty of the task. However, the high average $\pm$1 Acc19 (82.3%), low Distance (0.65), and strong Quadratic Weighted Kappa (88.1%) suggest that most disagreements between annotators and the unified labels were minor. For more detailed results on IAA, see Habash et al. (2025).
Input Model Acc19 $\pm$1 Acc19 Dist QWK
Word CamelBERT-msa 54.4% 68.7% 1.20 79.1%
MARBERTv2 53.3% 68.0% 1.20 79.1%
AraBERTv02 55.8% 69.2% 1.17 79.2%
AraBERTv2 51.6% 65.9% 1.32 76.3%
Lex CamelBERT-msa 48.3% 64.4% 1.34 77.1%
MARBERTv2 50.1% 64.9% 1.31 77.0%
AraBERTv02 48.8% 65.4% 1.30 78.5%
AraBERTv2 50.1% 65.4% 1.29 77.7%
D3Tok CamelBERT-msa 54.8% 68.2% 1.21 78.2%
MARBERTv2 54.0% 68.5% 1.20 78.9%
AraBERTv02 54.8% 68.1% 1.22 78.2%
AraBERTv2 56.6% 69.9% 1.14 80.0%
D3Lex CamelBERT-msa 51.1% 65.5% 1.29 78.0%
MARBERTv2 51.6% 65.7% 1.28 78.0%
AraBERTv02 53.3% 68.1% 1.24 78.2%
AraBERTv2 53.2% 67.1% 1.24 78.6%
Table 5: Results comparing different combinations of models and input variants on Barec Dev set. Bold are the best results on each metric.
Loss Acc19 $\pm$1 Acc19 Dist QWK
CE 56.6% 69.9% 1.14 80.0%
EMD 55.3% 70.3% 1.11 81.2%
OLL2 35.2% 70.3% 1.25 82.0%
OLL15 47.3% 71.1% 1.13 82.8%
OLL1 50.8% 71.5% 1.12 81.7%
OLL05 53.1% 68.8% 1.18 79.7%
SOFT2 55.8% 69.8% 1.15 80.0%
SOFT3 56.4% 69.9% 1.14 80.1%
SOFT4 56.4% 69.9% 1.15 79.6%
SOFT5 56.2% 69.5% 1.17 79.3%
Reg 43.1% 73.1% 1.13 84.0%
Table 6: Loss functions comparisons on Barec Dev set. We use AraBERTv2 model and D3Tok input with all loss function. Bold are the best results on each metric.
Input Loss Acc19 $\pm$1 Acc19 Dist QWK Acc7 Acc5 Acc3
Word CE 51.6% 65.9% 1.32 76.3% 61.6% 67.2% 74.0%
Lex CE 50.1% 65.4% 1.29 77.7% 60.6% 66.3% 74.9%
D3Tok CE 56.6% 69.9% 1.14 80.0% 65.9% 70.3% 76.5%
D3Lex CE 53.2% 67.1% 1.24 78.6% 63.6% 69.0% 75.3%
D3Tok EMD 55.3% 70.3% 1.11 81.2% 65.2% 70.0% 76.4%
D3Tok Reg 43.1% 73.1% 1.13 84.0% 61.1% 67.8% 75.9%
Average 46.9% 72.5% 1.11 83.4% 64.0% 70.3% 77.2%
Most Common 56.3% 70.0% 1.13 80.4% 66.3% 70.9% 76.9%
Oracle Combo 75.2% 87.4% 0.50 93.8% 83.2% 85.7% 89.1%
Table 7: Results comparing different loss function, ensemble methods, and oracle performance on Barec Dev set. Bold are the best results across individual models and across ensembles.
Input Loss Acc19 $\pm$1 Acc19 Dist QWK Acc7 Acc5 Acc3
Word CE 51.1% 65.1% 1.31 76.2% 60.7% 65.6% 72.2%
Lex CE 51.2% 66.2% 1.23 78.5% 61.1% 66.2% 74.4%
D3Tok CE 55.9% 70.0% 1.12 80.2% 65.1% 69.4% 75.2%
D3Lex CE 53.7% 67.9% 1.17 79.5% 63.8% 69.1% 74.8%
D3Tok EMD 54.9% 71.4% 1.02 83.7% 64.9% 69.0% 75.2%
D3Tok Reg 41.4% 73.5% 1.11 84.4% 59.4% 65.3% 72.8%
Average 46.0% 73.4% 1.06 84.5% 63.6% 69.4% 75.8%
Most Common 56.2% 70.4% 1.07 81.3% 65.9% 70.0% 75.6%
Oracle Combo 75.9% 87.8% 0.46 94.7% 83.5% 85.7% 88.9%
Table 8: Results comparing different loss function, ensemble methods, and oracle performance on Barec Test set. Bold are the best results across individual models and across ensembles.
6.2 Stage 1 Results
Table 5 presents the results of stage 1, where we evaluate different combinations of pretrained models and input variants using cross-entropy loss. Based on the all metrics, we observe that the AraBERTv02 and AraBERTv2 models generally achieve higher performance across multiple input variants. Among input variants, the Word and D3Tok representations tend to yield better results compared to Lex and D3Lex. Specifically, AraBERTv2 with the D3Tok input achieves the best scores in all metrics. Notably, AraBERTv2 is the only model that benefits from the D3Tok and D3Lex inputs compared to the Word input, showing an improvement across all metrics. We argue that this occurs because AraBERTv2 is the only model in this set that was pretrained on segmented data, making it more compatible with morphologically segmented input. These results suggest that both the choice of input variant and the pretrained model significantly impact performance. Based on all metrics, we select AraBERTv2 with the D3Tok input as the best-performing combination. In stage 2, we evaluate it with different loss functions. The confusion matrix for this model is available in the Appendix D.1.
6.3 Stage 2 Results
Table 6 presents the results of stage 2, where we use the best model from stage 1 to evaluate different loss functions. Among all the loss functions evaluated, Cross-Entropy (CE) achieves the highest exact accuracy (Acc19) at 56.6%, indicating that it performs best when predicting the exact readability level. In contrast, other loss functions show stronger performance on metrics that consider the ordinal nature of readability levels. Notably, Regression achieves the highest $\pm$1 accuracy at 73.1% and the best Quadratic Weighted Kappa (QWK) at 84.0%, suggesting it excels at predicting levels close to the gold label, despite being the worst in terms of exact accuracy. These findings support that loss functions designed for ordinal or continuous labels—such as EMD, OLL, and Regression—are more effective on evaluation metrics that reward proximity to the correct label, even if they underperform on strict accuracy. More results for other loss functions are in Appendix D.2.
6.4 Ensemble Results
Table 7 presents results from Stage 1, where AraBERTv2 is evaluated with four different input variants, and Stage 2, where it is trained using the two best-performing loss functions. It also includes results from two ensemble strategies applied across all six models to assess whether combining predictions can further improve performance. We also include an oracle combination, which represents an upper bound on performance. This allows us to estimate the maximum potential gain achievable through ensembling.
Ensemble
To further improve performance, we experiment with ensemble methods. We define the Average ensemble, where the final prediction is the rounded average of the levels predicted by the six models, and the Most Common ensemble, where the final prediction is the predicted levels’ mode. The results show that the Average ensemble performs better in terms of Distance, indicating that it tends to stay closer to the correct label. However, it struggles with exact accuracy (Acc), as averaging can blur distinctions between classes. On the other hand, the Most Common ensemble achieves higher Acc but can sometimes be misled by an incorrect majority, leading to greater deviation from the correct label.
Oracle
We also report an Oracle Combination, where we assume access to the best possible prediction from the six models for each sample. This serves as an upper bound on model performance. The Oracle results are significantly higher than those of individual models and are comparable to human annotators’ agreement with the unified labels (see section 6.1). This suggests that while individual models are still far from human-level performance, ensembling has the potential to push results closer to human agreement. More oracle combinations are provided in Appendix D.4. We also include more results on the impact of training granularity on readability level prediction in Appendix D.3 Finally, table 8 shows the results on the test set. We note that the trends observed in the development set persist in the test set, further validating our findings.
6.5 Error Analysis
To assess the errors in our best-performing model, we analyzed error patterns in the inter-annotator portion of the development (DEV) set. Each sentence in this subset had five human annotations, which we compared to the model’s prediction. We grouped sentences by the level of annotator agreement, from full agreement (5 out of 5 annotators) down to minimal agreement (1 out of 5). Full 5-way agreement accounts for 25% of the data. With each reduction in agreement – to 4, 3, 2, and finally 1 annotator – the cumulative coverage increases to 50%, 61%, 72%, and 87%, respectively. In other words, in 87% of the cases, the model prediction can be meaningfully compared to at least some level of human consensus. The remaining 13% fall outside this range. In 1% of these, the model’s prediction was within the span of human annotations but did not exactly match any of them. In 3%, the prediction was above the maximum annotation, and in 9%, it was below the minimum. We manually reviewed these out-of-range cases and found that the annotators were generally correct. We speculate that the model’s errors arise from limited training data, lack of contextual understanding, or insufficient modeling of linguistic features. For example, the obscure word \<عصامة> $\varsigma$$\hbar$ ‘tightly wound head dress’ may be misinterpreted as the feminine form of the proper name \<عصام> $\varsigma$SAm ‘Esam’, much like connecting \<كريم> krym ‘Kareem’ with \<كريمة> $\hbar$ ‘Kareema’. However, \<عصامة> $\varsigma$$\hbar$ is not a plausible proper name. This remains speculative, as our model is not inherently interpretable.
7 Conclusions and Future Work
This paper presented the Balanced Arabic Readability Evaluation Corpus (BAREC), a large-scale, finely annotated dataset for assessing Arabic text readability across 19 levels. With over 69K sentences and 1 million words, it is the largest Arabic corpus for readability assessment, covering diverse genres, topics, and audiences, to our knowledge. High inter-annotator agreement ensures reliable annotations. Through benchmarking various readability assessment techniques, we highlighted both the challenges and opportunities in Arabic readability modeling, demonstrating promising performance across different methods. Looking ahead, we plan to expand the corpus, enhancing its size and diversity to cover additional genres and topics. We also aim to add annotations related to vocabulary leveling and syntactic treebanks to study less-explored genres in syntax. Future work will include analyzing readability differences across genres and topics. Additionally, the tools we have developed will be integrated into a system to help children’s story writers target specific reading levels. The Barec dataset, its annotation guidelines, and benchmark results, will be made publicly available to support future research and educational applications in Arabic readability assessment.
Acknowledgments
The Barec project is supported by the Abu Dhabi Arabic Language Centre (ALC) / Department of Culture and Tourism, UAE. We acknowledge the support of the High Performance Computing Center at New York University Abu Dhabi. We are deeply grateful to our outstanding annotation team: Mirvat Dawi, Reem Faraj, Rita Raad, Sawsan Tannir, and Adel Wizani, Samar Zeino, and Zeina Zeino. Special thanks go to Abdallah Abushmaes, Karin Aghadjanian, and Omar Al Ayyoubi of the ALC for their continued support. We would also like to thank the Zayed University ZAI Arabic Language Research Center team, in particular Hamda Al-Hadhrami, Maha Fatha, and Metha Talhak, for their valuable contributions to typing materials for the project. We also acknowledge Ali Gomaa and his team for their additional support in this area. Finally, we thank our colleagues at the New York University Abu Dhabi Computational Approaches to Modeling Language (CAMeL) Lab, Muhammed Abu Odeh, Bashar Alhafni, Ossama Obeid, and Mostafa Saeed, as well as Nour Rabih (Mohamed bin Zayed University of Artificial Intelligence) for their helpful conversations and feedback.
Limitations
One notable limitation is the inherent subjectivity associated with readability assessment, which may introduce variability in annotation decisions despite our best efforts to maintain consistency. Additionally, the current version of the corpus may not fully capture the diverse linguistic landscape of the Arab world. Finally, while our methodology strives for inclusivity, there may be biases or gaps in the corpus due to factors such as selection bias in the source materials or limitations in the annotation process. We acknowledge that readability measures can be used with malicious intent to profile people; this is not our intention, and we discourage it.
Ethics Statement
All data used in the corpus curation process are sourced responsibly and legally. The annotation process is conducted with transparency and fairness, with multiple annotators involved to mitigate biases and ensure reliability. All annotators are paid fair wages for their contribution. The corpus and associated guidelines are made openly accessible to promote transparency, reproducibility, and collaboration in Arabic language research.
Appendix A Barec Annotation Guidelines Cheat Sheet and Examples
A.1 Arabic Original
A.2 English Translation
A.3 Annotation Examples
Representative examples of the 19 Barec readability levels, with English translations, and readability level reasoning. Underlining is used to highlight the main keys that determined the level. [Uncaptioned image]
Appendix B Barec Corpus Splits
B.1 Sentence-level splits across readability levels
B.2 Sentence-level splits across domains and readership groups
Appendix C Barec Corpus Details
C.1 Resources
We present the corpus sources in groups of their general intended purpose.
C.1.1 Education
Emarati Curriculum
The first five units of the UAE curriculum textbooks for the 12 grades in three subjects: Arabic language, social studies, Islamic studies Khalil et al. (2018).
ArabicMMLU
6,205 question and answer pairs from the ArabicMMLU benchmark dataset Koto et al. (2024).
Zayed Arabic-English Bilingual Undergraduate Corpus (ZAEBUC)
100 student-written articles from the Zayed University Arabic-English Bilingual Undergraduate Corpus Habash and Palfreyman (2022).
Arabic Learner Corpus (ALC)
16 L2 articles from the Arabic Learner Corpus (Alfaifi, 2015).
Basic Travel Expressions Corpus (BTEC)
20 documents from the MSA translation of the Basic Traveling Expression Corpus Eck and Hori (2005); Takezawa et al. (2007); Bouamor et al. (2018).
Collection of Children poems
Example of the included poems: My language sings (\<لغتي تغني>), and Poetry and news (\<أشعار وأخبار>) Al-Safadi (2005); Taha-Thomure (2007).
ChatGPT
To add more children’s materials, we ask Chatgpt to generate 200 sentences ranging from 2 to 4 words per sentence, 150 sentences ranging from 5 to 7 words per sentence and 100 sentences ranging from 8 to 10 words per sentence. [^5] https://chatgpt.com/ Not all sentences generated by ChatGPT were correct. We discarded some sentences that were flagged by the annotators. Table 9 shows the prompts and the percentage of discarded sentences for each prompt.
Table 9: ChatGPT Prompts. % Discarded is the percentage of discarded sentences due to grammatical errors.
C.1.2 Literature
Hindawi
A subset of 264 books extracted from the Hindawi Foundation website across different different genres. [^6] https://www.hindawi.org/books/categories/
Kalima
The first 500 words of 62 books from Kalima project. [^7] https://alc.ae/publications/kalima/
Green Library
58 manually typed books from the Green Library. [^8] https://archive.org/details/201409_201409
Arabian Nights
The openings and endings of the opening narrative and the first eight nights from the Arabian Nights Unknown (12th century). We extracted the text from an online forum. [^9] http://al-nada.eb2a.com/1000lela&lela/
Hayy ibn Yaqdhan
A subset of the philosophical novel and allegorical tale written by Ibn Tufail Tufail (1150). We extracted the text from the Hindawi Foundation website. [^10] https://www.hindawi.org/books/90463596/
Sara
The first 1000 words of Sara, a novel by Al-Akkad first published in 1938 Al-Akkad (1938). We extracted the text from the Hindawi Foundation website. [^11] https://www.hindawi.org/books/72707304/
The Suspended Odes (Odes)
The ten most celebrated poems from Pre-Islamic Arabia (\<المعلقات> Mu’allaqat). All texts were extracted from Wikipedia. [^12] https://ar.wikipedia.org/wiki/\<المعلقات¿
C.1.3 Media
Majed
10 manually typed editions of Majed magazine for children from 1983 to 2019.131313https://archive.org/details/majid_magazine
ReadMe++
The Arabic split of the ReadMe++ dataset Naous et al. (2024).
Spacetoon Songs
The opening songs of 53 animated children series from Spacetoon channel.
Subtitles
A subset of the Arabic side of the OpenSubtitles dataset Lison and Tiedemann (2016).
WikiNews
62 Arabic WikiNews articles covering politics, economics, health, science and technology, sports, arts, and culture Abdelali et al. (2016).
C.1.4 References
Wikipedia
A subset of 168 Arabic wikipedia articles covering Culture, Figures, Geography, History, Mathematics, Sciences, Society, Philosophy, Religions and Technologies. [^14] https://ar.wikipedia.org/
Constitutions
The first 2000 words of the Arabic constitutions from 16 Arabic speaking countries, collected from MCWC dataset El-Haj and Ezzini (2024).
UN
The Arabic translation of the Universal Declaration of Human Rights. [^15] https://www.un.org/ar/about-us/universal-declaration-of-human-rights
C.1.5 Religion
Old Testament
The first 20 chapters of the Book of Genesis Smith and Van Dyck (1865). [^16] https://www.arabicbible.com/
New Testament
[…]
Quran
The first three Surahs and the last 14 Surahs from the Holy Quran. We selected the text from the Quran Corpus Project Dukes et al. (2013). [^17] https://corpus.quran.com/
Hadith
The first 75 Hadiths from Sahih Bukhari al Bukhari (846). We selected the text from the LK Hadith Corpus [^18] https://github.com/ShathaTm/LK-Hadith-Corpus Altammami et al. (2019). Some datasets are chosen because they already have annotations available for other tasks. For example, dependency treebank annotations exist for Odes, Quran, Hadith, 1001, Hayy, OT, NT, Sara,WikiNews, ALC, BTEC, and ZAEBUC Habash et al. (2022a).
C.2 Domains
Arts & Humanities
The Arts and Humanities domain comprised the following subdomains.
  • Literature and Fiction: Encompasses novels, short stories, poetry, and other creative writing forms that emphasize narrative and artistic expression.
  • Religion and Philosophy: Contains religious texts, philosophical works, and related writings that explore spiritual beliefs, ethics, and metaphysical ideas.
  • Education and Academic Texts (on Arts and Humanities): Includes textbooks, scholarly articles, and educational materials that are often structured for learning and academic purposes.
  • General Knowledge and Encyclopedic Content (on Arts and Humanities): Covers reference materials such as encyclopedias, almanacs, and general knowledge articles that provide broad information on various topics.
  • News and Current Affairs (on Arts and Humanities): Includes newspapers, magazines, and online news sources that report on current events and issues affecting society.
Social Sciences
The Social Sciences domain comprised the following subdomains.
  • Business and Law: Encompasses legal texts, business strategies, financial reports, and corporate documentation relevant to professional and legal contexts.
  • Social Sciences and Humanities: Covers disciplines like sociology, anthropology, history, and cultural studies, which explore human society and culture.
  • Education and Academic Texts (on Social Sciences): Includes textbooks, scholarly articles, and educational materials that are often structured for learning and academic purposes.
  • General Knowledge and Encyclopedic Content (on Social Sciences): Covers reference materials such as encyclopedias, almanacs, and general knowledge articles that provide broad information on various topics.
  • News and Current Affairs (on Social Sciences): Includes newspapers, magazines, and online news sources that report on current events and issues affecting society.
STEM
The Science, Technology, Engineering and Mathematics domain comprised the following subdomains.
  • Science and Technology: Includes scientific research papers, technology articles, and technical manuals that focus on advancements and knowledge in science and tech fields.
  • Education and Academic Texts (on STEM): Includes textbooks, scholarly articles, and educational materials that are often structured for learning and academic purposes.
  • General Knowledge and Encyclopedic Content (on STEM): Covers reference materials such as encyclopedias, almanacs, and general knowledge articles that provide broad information on various topics.
  • News and Current Affairs (on STEM): Includes newspapers, magazines, and online news sources that report on current events and issues affecting society.
C.3 Readership Groups
Foundational
This level includes learners, typically up to 4th grade or age 10, who are building basic literacy skills, such as decoding words and understanding simple sentences.
Advanced
Refers to individuals with average adult reading abilities, capable of understanding a variety of texts with moderate complexity, handling everyday reading tasks with ease.
Specialized
Represents readers with advanced skills, typically starting in 9th grade or above in specialized topics, who can comprehend and engage with complex, domain-specific texts in specialized fields.
Table 10: Barec Corpus Details: the texts used to build the dataset, and the number of documents, sentences, and words extracted from each text.
Appendix D Additional Results
D.1 Confusion Matrix
Figure 3 shows the confusion matrix for the best-performing model from Stage 1: the AraBERTv2 model trained on D3Tok sentences with Cross-Entropy (CE) loss. The matrix uses F-scores to account for the unbalanced distribution of readability levels. The strong diagonal indicates a high rate of exact matches between predicted and gold labels. However, the model exhibits more disagreement at the higher, more difficult levels— likely due to the scarcity of training examples in those levels. Additionally, the model shows a tendency to under-estimate readability levels, favoring lower labels. This aligns with the patterns observed in the error analysis discussed in Section 6.5.
Figure 3: Confusion matrix of F-score across the different readability levels for the best model from stage 1.
D.2 All Loss Functions
Input Model Acc19 $\pm$1 Acc19 Dist QWK
Word SVM 36.2% 47.9% 2.03 53.4%
D3Tok SVM 37.2% 49.3% 1.92 56.5%
Word DecisionTree 27.2% 41.2% 2.50 44.2%
D3Tok DecisionTree 29.9% 44.2% 2.33 52.5%
D3Tok AraBERTv2
+CE 56.6% 69.9% 1.14 80.0%
+EMD 55.3% 70.3% 1.11 81.2%
+OLL2 35.2% 70.3% 1.25 82.0%
+OLL15 47.3% 71.1% 1.13 82.8%
+OLL1 50.8% 71.5% 1.12 81.7%
+OLL05 53.1% 68.8% 1.18 79.7%
+SOFT2 55.8% 69.8% 1.15 80.0%
+SOFT3 56.4% 69.9% 1.14 80.1%
+SOFT4 56.4% 69.9% 1.15 79.6%
+SOFT5 56.2% 69.5% 1.17 79.3%
+Reg 43.1% 73.1% 1.13 84.0%
Table 11: Loss functions comparisons on Barec Dev set. For SVM and Decision Tree classifiers, we used count vectorizer.
D.3 Impact of Training Granularity on Readability Level Prediction
To analyze the effect of training granularity on readability level prediction, we compare two approaches: (1) training on all 19 levels and then mapping predictions to lower levels (7, 5, or 3), and (2) training directly on the target granularity. Table 12 presents the results of this comparison. Overall, training on 19 levels and then mapping achieves slightly better performance across for 5-level and 3-level granularities compared to direct training. Moreover, the performance gap between the two approaches widens as the target granularity becomes coarser, suggesting that finer-grained supervision during training provides more informative learning signals, which translate into improved generalization when predictions are mapped into broader scales.
Train Gran Dev Gran Input Model Acc $\pm$1 Acc Dist QWK
19 7 D3Tok CE 65.9% 88.9% 0.51 79.9%
7 7 D3Tok CE 65.2% 89.5% 0.50 81.0%
19 5 D3Tok CE 70.3% 93.5% 0.37 78.3%
5 5 D3Tok CE 67.8% 93.7% 0.39 77.3%
19 3 D3Tok CE 76.5% 97.6% 0.26 74.7%
3 3 D3Tok CE 74.4% 96.9% 0.29 74.0%
Table 12: Comparison between training on 19 levels then mapping to the target granularity vs. training directly on the target granularity.
D.4 Ensembles & Oracles
CE CE CE CE EMD Reg Metrics
Word Lex D3Tok D3Lex D3Tok D3Tok Acc19 $\pm$1 Acc19 Dist QWK
51.6% 65.9% 1.32 76.3%
50.1% 65.4% 1.29 77.7%
56.6% 69.9% 1.14 80.0%
53.2% 67.1% 1.24 78.6%
55.3% 70.3% 1.11 81.2%
43.1% 73.1% 1.13 84.0%
Average 46.9% 72.5% 1.11 83.4%
Most Common 56.3% 70.0% 1.13 80.4%
Oracle Combinations
62.4% 76.6% 0.88 88.4%
63.5% 76.7% 0.89 87.7%
63.2% 76.6% 0.88 88.2%
63.3% 77.9% 0.83 89.2%
62.0% 80.7% 0.77 90.8%
69.5% 82.3% 0.67 91.4%
72.0% 84.5% 0.59 92.6%
73.6% 86.6% 0.53 93.4%
75.2% 87.4% 0.50 93.8%
Table 13: Comparison between individual models, ensembles and oracles on Barec Dev set.